OpenStudy (anonymous):

**MEDAL** I'm creating flashcards to help me study for the ASVAB. Help me with some formulas? A cardboard box has a length of 3 feet, height of 2½ feet, and depth of 2 feet. If the length and depth are doubled, by what percent does the volume of the box change? A. 200% B. 300% <-- C. 400% D. 600% Formula?

OpenStudy (anonymous):

Volume change= volume of the changed - volume of the original to get that into percent you multiply the resultant by (100/15)

OpenStudy (anonymous):

vi=15 vf= 60 change in v is 45 and then multiply it by 100/15

OpenStudy (anonymous):

Thanks again :)

OpenStudy (anonymous):

Real quick @M0j0jojo where did 60 and 15 come from? (I understand that 45 came from 60 - 15).

OpenStudy (anonymous):

The formula for volume of a the box is LWH so the original volume before the doubling is (3)(2.5)(2)=15 then the doubling happens and the dept and height are doubled so it becomes L2W2H= (3)(2x2.5)(2x2)=60 percent of increase = (V2 - V1)/V1 = (60 - 15)/15 = 45/15 = 3

OpenStudy (anonymous):

Ahh, that makes alot more sense! Fan for you, @M0j0jojo Thanks again :)

OpenStudy (anonymous):

No problem glad i could help :)